INVESTIGATION OF THE TORSIONAL FAR-INFRARED OVERTONES AND HOT BANDS OF AC-ETALDEHYDE AND VIBRATIONAL-TORSIONAL-ROTATIONAL INTERACTIONS WITH THE ν_{1}0 FUN-DAMENTAL BAND

\underline{I. KLEINER}, Laboratoire de Photophysique Moléculaire, Université Paris Sud, 91405 Orsay, France; N. MOAZZEN-AHMADI, Department of Physics, University of Lethbridge, Lethbridge, Alberta, T1K 3M4, Canada; A. R. W. McKELLAR, Stracie Institute for Molecular Science, National Research Council, Ottawa, Canada, K1AOR6; J. T. HOUGEN, Optical Physics Division, National Institute of Standards and Technology, Gaithersburg, MD 20899, USA.

The torsional overtone spectrum of CH_{3}CHO has been recorded with a high resolution (0.003 cm^{−1}) Far-Infrared Fourier Transform spectrometer, between 40 and 90 cm^{−1} and between 180 and 230 cm^{−1}. This spectrum completes the farinfrared spectrum recorded a few years ago from 80 to 180 cm^{−1} to study the fundamental torsional band v_{t} = 1–0 and the two first overtones v_{t} = 2–1 and 3–2^{a}. The spectral region between 40 and 90 cm^{−1} is very dense and allows us to assign many v_{t} = 2–1 lines and v_{t} = 3–2 lines. The region between 180 and 230 cm^{−1}, on the other hand, is less congested and allows us to search for weaker lines from the v_{t} = 3–1 and 4–2 overtone bands. Those overtone bands would correspond, in the free-rotor limit, to a change in the free-rotor quantum number m of ±3 and are expected to arise when considering intensity contribution from the second terms in the Fourier expansion of the dipole moment μ_{a}^{(3)} cos 3α and μ_{b}^{(3)} cos 3α for the A and E species and also from the first non-zero term of μ_{c}^{(3)} sin 3α for the E species. The spectrum of the ν_{10} band was recorded last year at high resolution with a Fourier Transform spectrometer in the region of 450 to 550 cm^{−1} region and partly assigned. A number of strong perturbations arising likely from interactions with the v_{t} = 4 torsional levels of the ground state were observed. As a first step to investigate doorway states which facilitate passage of energy from the small amplitude manifold to the torsional manifold, the goal of this study is to consider a two-by-two theoretical model involving both ν_{10} and the ground vibrational state and to carry out a simultaneous fit of the perturbed lines from v_{10}=1 and from vt=4 of the vibrational ground state.

aI. Kleiner, J. Hougen, J-U. Grabow, S. P. Belov, M. Yu. Tretyakov and J. Cosléou